OpenStudy (anonymous):

Which line is not perpendicular to the line 4x − 3y = 2?

OpenStudy (texaschic101):

4x - 3y = 2 first we need to find the slope of this line because a perpendicular slope will have the negative reciprocal. Do you know how to find the slope of this line ?

OpenStudy (texaschic101):

oops..a perpendicular line will have the negative reciprocal slope

OpenStudy (texaschic101):

Let me walk you through this...to find the slope we will put the equation in y = mx + b form where m is your slope. 4x - 3y = 2 --- subtract 4x from both sides -3y = -4x + 2 -- now divide both sides by -3 y = 4/3x - 2/3 the number in the m position (the slope) is 4/3 to find the perpendicular line we need the negative reciprocal of 4/3. To find the neg rec of a number, you " flip " the number and change the sign. So the negative reciprocal of 4/3 is -3/4. Therefore, a perpendicular line will contain the slope of -3/4. Any other slope is not perpendicular.

OpenStudy (texaschic101):

if your answer choices are in standard form of Ax + By = C, then you have to put them in y = mx + b form and find the slope....now remember, the slope has to be -3/4 or it is not perpendicular.

OpenStudy (texaschic101):

oops...I read the q wrong...you are looking for the one that is NOT perpendicular...so the one that does NOT have -3/4 as the slope is your answer

OpenStudy (anonymous):

Which line is not perpendicular to the line 4x − 3y = 2? 3x + 4y = 2 3x − 4y = 2 -3x − 4y = 5 3x + 4y = -8

OpenStudy (anonymous):

4x-3y=2

OpenStudy (anonymous):

can you help me

OpenStudy (texaschic101):

we have established that the perpendicular line will have -3/4 as the slope...so we need the one that does not have -3/4 as the slope. Put your answer choices in y = mx + b form. 3x + 4y = 2 4y = -3x + 2 y = -3/4x + 1/2 --- slope here is -3/4...it IS perpendicular 3x - 4y = 2 -4y = -3x + 2 y = 3/4x - 1/2 -- slope here is 3/4....NOT perpendicular -3x - 4y = 5 -4y = 3x + 5 y = -3/4x - 5/4 -- slope of -3/4...perpendicular 3x + 4y = -8 4y = -3x - 8 y = -3/4x - 2 -- slope is -3/4...perpendicular

OpenStudy (texaschic101):

do you have ANY questions at all ?

OpenStudy (anonymous):

omg thanks sooooo muchhhh !

OpenStudy (anonymous):

i totally understand now .

OpenStudy (texaschic101):

good to hear...it can be a bit confusing...lol

OpenStudy (anonymous):

lol yes!